# Label Distribution Learning using the Squared Neural Family on the Probability Simplex

### v1

@@ 14 Yeast_spoem 2,465 24 2
Random 0.1490$\pm$0.0098 $\bullet$ 0.6359$\pm$0.0247 $\bullet$ 1.1999$\pm$0.0448 $\bullet$ 0.1409$\pm$0.0136 $\bullet$ 0.9054$\pm$0.0093 $\bullet$ 0.7950$\pm$0.0089 $\bullet$
Kmeans 0.1456$\pm$0.0076 $\bullet$ 0.6242$\pm$0.0176 $\bullet$ 1.1815$\pm$0.0320 $\bullet$ 0.1361$\pm$0.0072 $\bullet$ 0.9079$\pm$0.0068 $\bullet$ 0.7981$\pm$0.0077 $\bullet$
CoreSet 0.1484$\pm$0.0164 $\bullet$ 0.6322$\pm$0.0383 $\bullet$ 1.1962$\pm$0.0720 $\bullet$ 0.1399$\pm$0.0179 $\bullet$ 0.9042$\pm$0.0158 $\bullet$ 0.7941$\pm$0.0171 $\bullet$
Graph Density 0.1428$\pm$0.0059 $\bullet$ 0.6175$\pm$0.0197 $\bullet$ 1.1671$\pm$0.0346 $\bullet$ 0.1326$\pm$0.0068 $\bullet$ 0.9108$\pm$0.0057 $\bullet$ 0.8013$\pm$0.0076 $\bullet$
Dirichlet 0.1472$\pm$0.0090 $\bullet$ 0.6282$\pm$0.0245 $\bullet$ 1.1899$\pm$0.0470 $\bullet$ 0.1376$\pm$0.0116 $\bullet$ 0.9064$\pm$0.0085 $\bullet$ 0.7959$\pm$0.0099 $\bullet$
SNEFY 0.1350$\pm$0.0030 0.5981$\pm$0.0134 1.1283$\pm$0.0254 0.1209$\pm$0.0049 0.9191$\pm$0.0030 0.8098$\pm$0.0042
@@
Table 3: The label distribution active learning performance on the Movie dataset.
@@
Label Distribution Active Learning
@@
In addition, we conduct experiments on label distribution active learning to evaluate SNEFY-LDL’s performance in uncertainty quantification. We choose the Movie dataset, randomly select 400 labeled training samples to form the initial labeled pool and take the remaining training samples as the unlabeled pool on each fold of cross-validation. The SNEFY-LDL model is first trained on the initial labeled pool. We then use different active learning strategies to pick up 200 informative samples from the unlabeled pool and query their labels. After augmenting the initial labeled pool with the 200 queried samples, we re-train the SNEFY-LDL model and evaluate its label distribution prediction performance on the test sets. Six different active learning strategies are compared:
@@
  • Random (Zhan et al., 2022) randomly selects 200 samples from the unlabeled pool.
@@
  • Kmeans (Zhdanov, 2019) selects samples close to the cluster centroids generated by the Kmeans clustering (MacKay, 2003) in feature space.
@@
  • CoreSet (Sener and Savarese, 2018) selects the $k$-center samples (Har-Peled, 2011) as representative unlabeled samples, which is a variant of Kmeans.
@@
  • Graph Density (Ebert et al., 2012) selects highly connected samples in the constructed KNN graph (Preparata and Shamos, 2012).
@@
  • Dirichlet models the distribution of label distributions using a Dirichlet distribution (Ng et al., 2011) centered at predicted label distributions and selects samples with the largest differential entropies (Cover, 1999).
@@
  • SNEFY-LDL uses importance sampling (Kloek and Van Dijk, 1978) to estimate differential entropies of the conditional distributions modeled by SNEFY-LDL and picks up samples having the largest differential entropies.
Among the six active learning strategies, Kmeans, CoreSet and Graph Density are representativeness based methods, which rely on only sample features by characterizing samples’ geometric properties in feature space, while Dirichlet and SNEFY-LDL are the uncertainty based methods we have contrived, which leverage the predictions of the initially trained SNEFY-LDL model.
Table 3 shows the performance of different active learning strategies, where the best performer is highlighted by boldface and $\bullet$ indicates that the annotated performer is inferior to the best performer at 0.05 significance level. As is shown in Table 3, SNEFY-LDL consistently achieves the best performance in terms of all metrics. By accurately evaluating label distribution prediction uncertainties, SNEFY-LDL can pick up more informative unlabeled samples than the naive uncertainty quantification strategy, Dirichlet, as well as the representativeness based active learning strategies, Kmeans, CoreSet and Graph Density, which are even sometimes inferior to the Random strategy.
@@
SA-BFGS Average 0.1178$\pm$0.0020$\bullet$ 0.3743$\pm$0.0068$\bullet$ 0.7948$\pm$0.0163$\bullet$ 0.0641$\pm$0.0022$\bullet$ 0.9370$\pm$0.0020$\bullet$ 0.8575$\pm$0.0028$\bullet$
@@
DF-LDL Average 0.1203$\pm$0.0017$\bullet$ 0.3762$\pm$0.0059$\bullet$ 0.8040$\pm$0.0146$\bullet$ 0.0657$\pm$0.0019$\bullet$ 0.9353$\pm$0.0018$\bullet$ 0.8557$\pm$0.0025$\bullet$
@@
LDL-SCL Average 0.1256$\pm$0.0020$\bullet$ 0.3828$\pm$0.0047$\bullet$ 0.8260$\pm$0.0118$\bullet$ 0.0699$\pm$0.0021$\bullet$ 0.9315$\pm$0.0018$\bullet$ 0.8519$\pm$0.0020$\bullet$
@@
LDL-LRR Average 0.1269$\pm$0.0021$\bullet$ 0.3966$\pm$0.0052$\bullet$ 0.8478$\pm$0.0131$\bullet$ 0.0730$\pm$0.0022$\bullet$ 0.9285$\pm$0.0019$\bullet$ 0.8478$\pm$0.0023$\bullet$
@@
Label Distribution Ensemble Learning
We also conduct experiments on ensemble learning to further verify SNEFY-LDL’s ability in uncertainty quantification, by using Bagging (Breiman, 1996) as an exemplary ensemble learning paradigm. We choose the SBU_3DFE dataset, randomly select 50 samples from the training split of each fold of cross-validation for 25 rounds, train 25 base LDL learners with the selected samples, and evaluate the label distribution prediction performance of ensembled LDL models on the test sets. Four competitive LDL models are employed as base learners: SA-BFGS, DF-LDL, LDL-SCL and LDL-LRR, and two strategies are adopted to ensemble base learner predictions: 1) Average: aggregate the 25 base learner predictions with the uniform weight $1/25$, and 2) Weighted: weight each base learner prediction in proportion to its SNEFY-LDL conditional likelihood in an instance-wise manner.
Table 4 compares the two different ensemble learning strategies, where the best strategy is highlighted by boldface and $\bullet$ indicates that the annotated strategy is inferior to the best strategy at 0.05 significance level. From Table 4, we can find that the Weighted strategy is significantly better than Average in terms of all metrics. This implies that SNEFY-LDL provides an effective mechanism to quantify the prediction reliability/uncertainties of base learners.
Parameter Sensitivity Study
We also conduct additional experiments to investigate SNEFY-LDL’s sensitivity with regard to its four hyperparameters: $n$ and $m$, as well as the batch size and epoch number used for training. The experiments demonstrate that the performance of SNEFY-LDL remains relatively stable when the hyperparameters vary in proper ranges. Please check the Appendix for more details.
@@ 7 Conclusion
In this paper, we propose a novel LDL paradigm: estimate the distribution of label distribution vectors in the probability simplex, which brings a bird’s-eye view on the relative significance of all possible label distributions. Through uncovering the underlying relationship between SNEFY and LDL, we develop the SNEFY-LDL model that can provide a tractable formulation of the conditional distribution of label distribution vectors, enjoying great expressivity and high computational efficiency. SNEFY-LDL admits closed-form expressions for the distribution’s mean, variance and covariance that can be computed quickly, making SNEFY-LDL able to provide real-time responses to the downstream tasks and real-world applications. Experiments on label distribution prediction, as well as label distribution active learning and ensemble learning demonstrate the great utility of SNEFY-LDL for label distribution modeling.
@@ Appendix A Theorem Proofs
As $\bm{W}_{1}>-1/2$ elementwise, $w_{1il}+w_{1jl}+1>0$. Assuming $\bm{\ell}$ follows a Dirichlet distribution (Ng, Tian, and Tang 2011) with parameters $\bm{\alpha}=(\alpha_{1},\alpha_{2},\cdots,\alpha_{L})$, where $\alpha_{l}=w_{1il}+w_{1jl}+1>0$, its probability density, $\mathrm{P}_{\mathrm{Dir}}(d\bm{\ell})/{d\bm{\ell}}$, is in the form:
@@
$\int_{\Delta^{L-1}}\prod_{l=1}^{L}(\bm{\ell}^{y_{l}})^{\alpha_{l}-1}d\bm{\ell}=\mathrm{B}(\alpha).$
@@
Assuming the label distribution vector $\bm{\ell}$ follows the SNEFY conditional distribution $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$ in Eq. (7) with the kernel function $\bm{k}_{\sigma,\bm{t}_{1},\bm{t}_{2},\mu_{1}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{x})$ given in Eq. (10), under the setting that $\bm{t}_{1}(\bm{\ell})=(\log\bm{\ell}^{y_{1}},\log\bm{\ell}^{y_{2}},\cdots,\log\bm{\ell}^{y_{L}})$, $\sigma=\exp$, and $\mu_{1}(d\bm{\ell})=d\bm{\ell}$, as well as the constraint that $\bm{W}_{1}>-1/2$ elementwise, for the affiliation probability to the $r$th label $y_{r}\in\mathcal{Y}$, $\bm{\ell}^{y_{r}}$, we have its conditional mean $\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]$ as
@@
For any fixed function $\varphi(\bm{\ell})$ of $\bm{\ell}$, its expectation with regard to the conditional SNEFY distribution in Eq. (7), $\mathrm{E}[\varphi(\bm{\ell})|\bm{x}]$, can be computed as
@@
$=\exp(\bm{w}_{2i}^{\top}\bm{t}_{2}(\bm{x})+\bm{w}_{2j}^{\top}\bm{t}_{2}(\bm{x})+b_{i}+b_{j})\cdot\int_{\Delta^{L-1}}\varphi(\bm{\ell})\prod_{l=1}^{L}(\bm{\ell}^{y_{l}})^{w_{1il}+w_{1jl}}d\bm{\ell}$
@@
$\bm{\phi}_{\bm{t}_{2}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{x})=$ $\exp(\bm{w}_{2i}^{\top}\bm{t}_{2}(\bm{x})+\bm{w}_{2j}^{\top}\bm{t}_{2}(\bm{x})+b_{i}+b_{j})$
$\cdot$ $\frac{\prod_{l=1}^{L}\Gamma(1+w_{1il}+w_{1jl})}{\Gamma\big{(}L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})\big{)}}\mathrm{E}_{\mathrm{Dir}}[\varphi(\bm{\ell});\bm{w}_{1i},\bm{w}_{1j}]$
@@
$\mathrm{E}_{\mathrm{Dir}}[\bm{\ell}^{y_{r}};\bm{w}_{1i},\bm{w}_{1j}]=\frac{1+w_{1ir}+w_{1jr}}{L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})}.$
$\mathrm{E}_{\mathrm{Dir}}[(\bm{\ell}^{y_{r}})^{2};\bm{w}_{1i},\bm{w}_{1j}]=\frac{(1+w_{1ir}+w_{1jr})(2+w_{1ir}+w_{1jr})}{[L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})][1+L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})]}.$

### v2

@@ 6.1 Conformal Prediction
Random 0.1490$\pm$0.0098 0.6359$\pm$0.0247 1.1999$\pm$0.0448 0.1409$\pm$0.0136 0.9054$\pm$0.0093 0.7950$\pm$0.0089
Kmeans 0.1456$\pm$0.0076 0.6242$\pm$0.0176 1.1815$\pm$0.0320 0.1361$\pm$0.0072 0.9079$\pm$0.0068 0.7981$\pm$0.0077
CoreSet 0.1484$\pm$0.0164 0.6322$\pm$0.0383 1.1962$\pm$0.0720 0.1399$\pm$0.0179 0.9042$\pm$0.0158 0.7941$\pm$0.0171
Graph Density 0.1428$\pm$0.0059 0.6175$\pm$0.0197 1.1671$\pm$0.0346 0.1326$\pm$0.0068 0.9108$\pm$0.0057 0.8013$\pm$0.0076
Dirichlet 0.1472$\pm$0.0090 0.6282$\pm$0.0245 1.1899$\pm$0.0470 0.1376$\pm$0.0116 0.9064$\pm$0.0085 0.7959$\pm$0.0099
SNEFY-LDL 0.1350$\pm$0.0030 0.5981$\pm$0.0134 1.1283$\pm$0.0254 0.1209$\pm$0.0049 0.9191$\pm$0.0030 0.8098$\pm$0.0042
@@
Table 4: The label distribution active learning performance on the Motive dataset.
Random 0.3591$\pm$0.0197 2.4871$\pm$0.0280 6.9131$\pm$0.1456 0.9835$\pm$0.0766 0.6554$\pm$0.0411 0.4528$\pm$0.0336
Kmeans 0.3690$\pm$0.0163 2.4929$\pm$0.0200 6.9407$\pm$0.0808 1.0498$\pm$ 0.1078 0.6318$\pm$0.0375 0.4361$\pm$0.0228
CoreSet 0.3629$\pm$0.0193 2.4842$\pm$0.0233 6.8882$\pm$0.1089 1.0102$\pm$0.1078 0.6442$\pm$0.0489 0.4485$\pm$0.0340
Graph Density 0.3700$\pm$0.0242 2.4958$\pm$0.0277 6.9643$\pm$0.1394 1.0471$\pm$0.0991 0.6230$\pm$0.0453 0.4294$\pm$0.0321
Dirichlet 0.3720$\pm$0.0213 2.4951$\pm$0.0226 6.9502$\pm$0.0977 1.0628$\pm$0.0909 0.6249$\pm$0.0336 0.4326$\pm$0.0192
SNEFY-LDL 0.3474$\pm$0.0160 2.4807$\pm$0.0194 6.8755$\pm$0.0845 0.9244$\pm$0.0648 0.6819$\pm$0.0285 0.4718$\pm$0.0203
@@
Table 5: The label distribution active learning performance on the Natural_Scene dataset.
As a baseline, we extend the competitive SA-BFGS [Geng, 2016] algorithm by modeling the distribution of (rather than point-estimating) the label distribution vectors. We model the distribution using the Dirichlet distribution centered at the point prediction. Following the same routine of conformal prediction, we can also construct calibrated confidence intervals for label affiliation probabilities given new samples. Following Angelopoulos and Bates [2021], we use the Feature-Stratified Coverage (FSC) metric to evaluate the adaptivity of the constructed confidence intervals in Eq.(19), which categorizes the test samples into different groups by dividing the numeric values at the first feature dimension into different bins, then compute the coverage rate of the confidence intervals in each group, and picks up the lowest group-level coverage rate as the final metric value. We select the Natural_Scene and SBU_3DFE datasets with a fair number of examples and labels to test the performance of conformal prediction. We randomly split each dataset into the training, calibration and test sets according to the ratio of 50%/25%/25% for 100 times and report the average FSC scores with bin size equal to 2, 4 and 8. In the experiment, we aim to construct 90% level confidence intervals by setting $0.11/k^{2}=$ 0.1, which means that FSC scores closer to 0.9 indicates better conformal prediction performance.
Tables 2-3 compare the conformal prediction performance measured by the FSC metric on the Natural_Scene and SBU_3DFE datasets. For each comparison between SNEFY-LDL and Dirichlet, the better performer is highlighted by boldface. From the tables, we can find that the proposed SNEFY-LDL model outperforms the Dirichlet baseline in most cases. SNEFY-LDL constructs a multimodal distribution to model the distribution of label distribution vectors on the probability simplex space, which is more flexible than the unimodal Dirichlet, and contributes to the label distribution confidence intervals with greater adaptivity.
6.2 Active Learning
@@
To further evaluate SNEFY-LDL’s performance in uncertainty quantification, we choose the Movie and Natural_Scene datasets to conduct the active learning experiments. We first randomly split the two datasets into the training and test sets according to the ratio of 90%/10% for ten times. For each training-test set split, we randomly select 400 labeled training samples to form the initial labeled pool and take the remaining training samples as the unlabeled pool. We first train a SNEFY-LDL model with the labeled samples in the initial labeled pool. We then use different active learning strategies to pick up 100 informative samples from the unlabeled pool and query their labels. After augmenting the initial labeled pool with the 100 queried samples, we re-train the SNEFY-LDL model and evaluate the performance of its label distribution predictions produced by the closed-form conditional mean in Eq. (10). Following Geng [2016], the label distribution prediction performance is evaluated by the following six metrics: Chebyshev distance (Cheby), Clark distance (Clark), Canberra metric (Canb), Kullback-Leibler divergence (KL), Cosine coefficient (Cos) and Intersection (Inter). The average scores on the ten random training-test splits are reported. Six different active learning strategies are compared:
@@
  • Random [Zhan et al., 2022] randomly selects 100 samples from the unlabeled pool.
@@
  • Kmeans [Zhdanov, 2019] selects samples close to the cluster centroids generated by the Kmeans clustering [MacKay, 2003] in feature space.
@@
  • CoreSet [Sener and Savarese, 2018] selects the $k$-center samples [Har-Peled, 2011] as representative unlabeled samples, which is a variant of Kmeans.
@@
  • Graph Density [Ebert et al., 2012] selects highly connected samples in the constructed KNN graph [Preparata and Shamos, 2012].
@@
  • Dirichlet models the distribution of label distributions using a Dirichlet distribution [Ng et al., 2011] centered at predicted label distributions and selects samples with the largest differential entropy scores [Cover, 1999].
@@
  • SNEFY-LDL uses importance sampling [Kloek and Van Dijk, 1978] to estimate the differential entropy values of the conditional distributions modeled by SNEFY-LDL and picks up samples that have the largest differential entropy scores.
@@
SA-BFGS Average 0.1178$\pm$0.0020 0.3743$\pm$0.0068 0.7948$\pm$0.0163 0.0641$\pm$0.0022 0.9370$\pm$0.0020 0.8575$\pm$0.0028
@@
DF-LDL Average 0.1203$\pm$0.0017 0.3762$\pm$0.0059 0.8040$\pm$0.0146 0.0657$\pm$0.0019 0.9353$\pm$0.0018 0.8557$\pm$0.0025
@@
LDL-SCL Average 0.1256$\pm$0.0020 0.3828$\pm$0.0047 0.8260$\pm$0.0118 0.0699$\pm$0.0021 0.9315$\pm$0.0018 0.8519$\pm$0.0020
@@
LDL-LRR Average 0.1269$\pm$0.0021 0.3966$\pm$0.0052 0.8478$\pm$0.0131 0.0730$\pm$0.0022 0.9285$\pm$0.0019 0.8478$\pm$0.0023
@@
Base Learner Bagging Cheby $\downarrow$ Clarky $\downarrow$ Canb $\downarrow$ KL $\downarrow$ Cos $\uparrow$ Inter $\uparrow$
SA-BFGS Average 0.0889$\pm$0.0085 0.3180$\pm$0.0197 0.6529$\pm$0.0461 0.0406$\pm$0.0058 0.9613$\pm$0.0057 0.8890$\pm$0.0086
Weighted 0.0842$\pm$0.0084 0.3118$\pm$0.0200 0.6390$\pm$0.0466 0.0385$\pm$0.0056 0.9636$\pm$0.0056 0.8923$\pm$0.0089
DF-LDL Average 0.0951$\pm$0.0093 0.3385$\pm$0.0253 0.6958$\pm$0.0611 0.0456$\pm$0.0078 0.9566$\pm$0.0072 0.8818$\pm$0.0110
Weighted 0.0881$\pm$0.0092 0.3189$\pm$0.0258 0.6525$\pm$0.0617 0.0408$\pm$0.0074 0.9612$\pm$0.0069 0.8895$\pm$0.0110
LDL-SCL Average 0.0911$\pm$0.0090 0.3249$\pm$0.0219 0.6746$\pm$0.0493 0.0424$\pm$0.0063 0.9596$\pm$0.0062 0.8854$\pm$0.0092
Weighted 0.0865$\pm$0.0095 0.3153$\pm$0.0231 0.6509$\pm$0.0540 0.0400$\pm$0.0064 0.9621$\pm$0.0064 0.8898$\pm$0.0100
LDL-LRR Average 0.0888$\pm$0.0090 0.3189$\pm$0.0209 0.6538$\pm$0.0473 0.0408$\pm$0.0065 0.9612$\pm$0.0063 0.8890$\pm$0.0090
Weighted 0.0846$\pm$0.0089 0.3124$\pm$0.0197 0.6392$\pm$0.0444 0.0387$\pm$0.0062 0.9634$\pm$0.0063 0.8922$\pm$0.0087
Table 7: The label distribution ensemble learning performance on the SJAFFE dataset.
Tables 4-5 compare the performance of different active learning strategies, where the best performer is highlighted by boldface. As is shown in the tables, SNEFY-LDL consistently achieves the best performance in terms of all metrics. By accurately evaluating label distribution prediction uncertainties, SNEFY-LDL can pick up more informative unlabeled samples than the naive uncertainty quantification strategy, Dirichlet, as well as the representativeness based active learning strategies, Kmeans, CoreSet and Graph Density, which are even sometimes inferior to the Random strategy.
6.3 Ensemble Learning
We also conduct experiments on ensemble learning to further verify SNEFY-LDL’s ability in uncertainty quantification, with the expectation that reliable base learners can be identified by the SNEFY-LDL probability modeling. Bagging [Breiman, 1996] is adopted as an exemplary ensemble learning paradigm. We choose the SEU_3DFE and SJAFFE datasets, and randomly split them into training and test sets according to the ratio of 90%/10%. For each training-test set split, we randomly select 50 samples from the training set for 25 rounds, train 25 base LDL learners with the selected samples, and evaluate the label distribution prediction performance of the ensembled LDL model on the test set. Four competitive LDL models are employed as base learners: SA-BFGS [Geng, 2016], DF-LDL [González et al., 2021a], LDL-SCL [Jia et al., 2021] and LDL-LRR [Jia et al., 2023a], and two strategies are adopted to ensemble base learner predictions: 1) Average: aggregate the 25 base learner predictions with the uniform weight $1/25$, and 2) Weighted: weight each base learner prediction in proportion to its corresponding SNEFY-LDL probability density conditioned on each test sample in an instance-wise manner.
Tables 6-7 compare the two different ensemble learning strategies, where the best strategy is highlighted by boldface. From the tables, we find that the Weighted strategy is significantly better than Average in terms of all metrics. This implies that SNEFY-LDL provides an effective mechanism to quantify the reliability of base learners’ label distribution predictions so that the reliable base learners are highlighted to contribute to a better ensemble learning performance.
@@ 7 Conclusion
We propose a novel LDL paradigm: estimate the distribution of label distribution vectors on the probability simplex, which brings a bird’s-eye view on the relative significance of all possible label distributions. By uncovering the underlying relationship between SNEFY and LDL, we develop the SNEFY-LDL model that can provide a tractable formulation of the conditional distribution of label distribution vectors, enjoying great expressivity and high computational efficiency. SNEFY-LDL admits closed-form expressions for the distribution’s mean, variance and covariance, making SNEFY-LDL able to provide real-time responses in real-world applications. Experiments on conformal prediction, active learning and ensemble learning demonstrate the great utility of SNEFY-LDL for uncertainty-aware applications.
@@ Appendix A Theorem Proofs
$\int_{\Delta^{L-1}}\prod_{l=1}^{L}(\bm{\ell}^{y_{l}})^{\alpha_{l}-1}d\bm{\ell}=\mathrm{B}(\bm{\alpha}).$
@@
Assuming the label distribution vector $\bm{\ell}$ follows the SNEFY conditional distribution $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$ in Eq. (6) with the kernel function $\bm{k}_{\sigma,\bm{t}_{1},\bm{t}_{2},\mu_{1}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{x})$ given in Eq. (9), under the setting that $\bm{t}_{1}(\bm{\ell})=(\log\bm{\ell}^{y_{1}},\log\bm{\ell}^{y_{2}},\cdots,\log\bm{\ell}^{y_{L}})$, $\sigma=\exp$, and $\mu_{1}(d\bm{\ell})=d\bm{\ell}$, as well as the constraint that $\bm{W}_{1}>-1/2$ elementwise, for the $r$th label’s composition ratio, $\bm{\ell}^{y_{r}}$, we have its conditional mean $\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]$ as
@@
For any fixed scalar function $\varphi(\bm{\ell})$ of $\bm{\ell}$ with $\varphi(\cdot):\Delta^{L-1}\rightarrow\mathbb{R}$, its expectation with regard to the conditional SNEFY distribution in Eq. (6), $\mathrm{E}[\varphi(\bm{\ell})|\bm{x}]$, can be computed as
@@
$=\exp(\bm{w}_{2i}^{\top}\bm{t}_{2}(\bm{x})+\bm{w}_{2j}^{\top}\bm{t}_{2}(\bm{x})+b_{i}+b_{j})\cdot\int_{\Delta^{L-1}}\varphi(\bm{\ell})\prod_{l=1}^{L}(\bm{\ell}^{y_{l}})^{w_{1il}+w_{1jl}}d\bm{\ell}.$
@@
$\bm{\phi}_{\bm{t}_{2}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{x})=$ $\exp(\bm{w}_{2i}^{\top}\bm{t}_{2}(\bm{x})+\bm{w}_{2j}^{\top}\bm{t}_{2}(\bm{x})+b_{i}+b_{j})\cdot\frac{\prod_{l=1}^{L}\Gamma(1+w_{1il}+w_{1jl})}{\Gamma\big{(}L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})\big{)}}\mathrm{E}_{\mathrm{Dir}}[\varphi(\bm{\ell});\bm{w}_{1i},\bm{w}_{1j}]$
@@
$\mathrm{E}_{\mathrm{Dir}}[\bm{\ell}^{y_{r}};\bm{w}_{1i},\bm{w}_{1j}]=\frac{1+w_{1ir}+w_{1jr}}{L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})},$
$\mathrm{E}_{\mathrm{Dir}}[(\bm{\ell}^{y_{r}})^{2};\bm{w}_{1i},\bm{w}_{1j}]=\frac{(1+w_{1ir}+w_{1jr})(2+w_{1ir}+w_{1jr})}{[L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})][1+L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})]},$